OpenStudy (anonymous):

If the line passing through the points (a, 1) and (2, 6) is parallel to the line passing through the points (0, 7) and (a + 2, 1), what is the value of a?

OpenStudy (anonymous):

Parallel lines=same slope \[\frac{6-1}{2-a}=\frac{1-7}{a+2}\] \[5(a+2)=-6(2-a)\] \[5a+10=-12+6a, a=22\]

OpenStudy (anonymous):

thank you!!